# Unsupervised Morphological Tree Tokenizer

### v1

@@ Abstract
As a cornerstone in language modeling, tokenization involves segmenting text inputs into pre-defined atomic units. Conventional statistical tokenizers often disrupt constituent boundaries within words, thereby corrupting semantic information. To address this drawback, we introduce morphological structure guidance to tokenization and propose a deep model to induce character-level structures of words. Specifically, the deep model jointly encodes internal structures and representations of words with a mechanism named MorphOverriding to ensure the indecomposability of morphemes. By training the model with self-supervised objectives, our method is capable of inducing character-level structures that align with morphological rules without annotated training data. Based on the induced structures, our algorithm tokenizes words through vocabulary matching in a top-down manner. Empirical results indicate that the proposed method effectively retains complete morphemes and outperforms widely adopted methods such as BPE and WordPiece on both morphological segmentation tasks and language modeling tasks. The code will be released later.
@@ 1 Introduction
Inspired by linguistic theories that words have internal structures Selkirk (1982); Marvin (2002); Cotterell and Schütze (2015), we seek to tokenize words based on their morphological structures as shown in Figure 1(bottom), aiming to enhance the alignment between tokenization and the morphological intricacies of language. Specifically, we consider an unsupervised approach to induce character-level morphological structures within words, which circumvents the need for annotated morphological data that are unavailable for many languages. Our approach draws inspiration from syntactic composition models Maillard et al. (2017), where a sentence is encoded as a weighted sum over all composed root representations of its underlying binary parse trees via dynamic programming. Instead of composing a sentence from words, we apply composition models on characters in a word to induce its morphological parse tree. To train the composition model, we propose two self-supervised objectives akin to next token prediction and span prediction that effectively leverage both contextual information at the sentence level and semantic information at the subword level. Thus the model can learn to assign higher probabilities to morphological constituents of a word and induce the underlying morphological parse tree.
@@
In our experiments, we train TreeTok and baselines on the Wikitext-103 corpus McClosky et al. (2006) and assess their performance on morphological segmentation tasks and language modeling tasks. Evaluation results indicate that TreeTok consistently outperforms BPE and WordPiece across all the tasks.
@@
We propose a composition model with MorphOverriding for modeling the compositionality within words.
We propose effective self-supervised objectives to train the composition model to induce morphological structures of words highly consistent with gold morphological segmentations.
We propose a novel tokenization algorithm based on induced morphological parse trees.
@@ 2 Related Work
Subword tokenization, with typical methods such as BPE Sennrich et al. (2016) and WordPiece Schuster and Nakajima (2012), has become customary in most NLP fields. BPE builds its vocabulary by repeatedly merging the most frequent subword unit pairs, whereas WordPiece selects pairs using the highest mutual information. During tokenization, both apply these learned merge operations in the same order to new text initialized with characters. Unigram Kudo (2018), another popular tokenizer, builds its vocabulary in the opposite direction: it starts with a large set of potential subwords and prunes them based on delta entropy in a unigram language model.
Our tokenizer aims to build upon the advantages of these effective statistical tokenizers and augment them with unsupervised induced tree structures.
@@
In the line of work on unsupervised morphological segmentation, the most well-known model is Morfessor Creutz and Lagus (2002), along with its multiple variants Creutz and Lagus (2005); Grönroos et al. (2014, 2020). In Morfessor, an online search algorithm is utilized to apply a hierarchical word splitting strategy with a Minimum Description Length (MDL) Rissanen (1989) cost function. However, although some studies Ataman and Federico (2018); Hou et al. (2023) find unsupervised morphologically motivated segmentation can improve data-driven tokenizers, most other studies Machácek et al. (2018); Domingo et al. (2019); Sälevä and Lignos (2021) find no reliable improvement of such methods over BPE. According to Gallé (2019), the effectiveness of BPE lies in its superior compression capability. A more detailed discussion can be found in Mielke et al. (2021).
Some other studies try to model morphological structures using Bayesian PCFGs Johnson et al. (2007) or a non-parametric Bayesian generalization of PCFGs Johnson et al. (2006). However, they are pure statistical models and do not utilize modern neural methodologies.
Our method differs from previous unsupervised morphological methods in our character-based structures, thereby possessing the superior compression capability of BPE. Meanwhile, our method leverages modern neural methodologies to better utilize contextual and intra-word semantic information.
@@
In this work, we utilize a composition model to induce morphological structures. Composition models jointly learn representations and structures of a symbol sequence by transforming text encoding into a combinatorial optimization problem. Maillard et al. (2017) proposes a CKY-like Cocke (1969); Kasami (1966); Younger (1967) encoder, in which each constituent is represented as a weighted average of the set of composed representations computed from different splits of the constituent. Drozdov et al. (2019) proposes a deep inside-outside encoder Baker (1979); Lari and Young (1990), enabling the encoder to learn underlying structures via an auto-encoding objective. Recently, a series of studies Hu et al. (2024a, b) have been conducted to reduce the deep inside-outside encoder complexity from cubic to linear, on which our work is based 111https://github.com/ant-research/StructuredLM_RTDT.
@@ 3 Methodology
Given a word where is the -th character, we aim to parse it into a binary tree and then tokenize it via top-down vocabulary matching. The parser is a deep composition model capable of jointly modeling the internal structures and representations of words and is trained by optimizing self-supervised objectives. In the following sections, we sequentially introduce the composition model, training objectives, and the tree-based tokenization algorithm.
@@ 3.2 Training Objectives
The overall loss for training the composition model is the summation of an auto-encoding loss and an auto-regression loss . The auto-encoding loss is based on predicting each character or morpheme from the rest of a word, leveraging intra-word information. The auto-regression loss is based on next token prediction that leverages contextual information to disambiguate the underlying structures of a word. Under these objectives, the composition model learns to assign proper scores to each split point of a subword.
Auto-encoding Loss.
Auto-encoding is a common practice of training a composition model. For our character-level composition model, we try to predict each character based on its neighboring context representations and Hu et al. (2021). However, the auto-encoding objective turns out to be empirically ineffective when training our model probably because unlike word-level auto-encoding that requires selecting from tens of thousands of words in a vocabulary, here we only need to select from tens of characters, which is much less challenging. To enhance learning efficacy, we propose predicting both individual characters and morphemes in the vocabulary . For instance, given the word windsurf, we mask out wind and let the model uncover the masked morpheme based on the visible part _surf. Analogous to the inside pass, the outside pass computes each outside representation in a top-down manner based on context information outside span , whose details are described in Appendix A.3. we use to predict each subword that belongs to :
where is the total number of subwords belonging to the vocabulary 222Note that multiple subwords may be mutually exclusive, such as “asking” with “ask” and “king”. Intuitively, it is more reasonable to predict a constituent than a distituent from its context. Therefore, we assign a constituency weight to each subword in the objective, as detailed in Appendix A.4.
Auto-regression Loss.
@@ 3.3 Tokenization
Given a constructed vocabulary, whose details are described later, we parse each word into a morphological tree and segment it via a top-down matching approach, as illustrated in Figure 1(bottom). Specifically, during the top-down traversal of a parse tree, we retain a subword and backtrack if the subword matches an entry in the vocabulary. Note that unsupervised structural learning is often imperfect, causing erroneous tokenization. For instance, an incorrect parse tree [[[book]e]d] may yield tokens book e d where e d should be merged. To address this issue, we propose a post-processing step to deal with mergeable pairs of segmented tokens. Specifically, we define the information entropy of a token as , where is the frequency of in the entire corpus. Therefore, the entropy of a certain merge is the sum of the information entropies of all tokens. We find the optimal merge by searching for the one with minimal information entropy among all potential merges via dynamic programming. Detailed pseudo-code can be found in Appendix A.1.
@@
Tokenize the corpus via the top-down matching according to the current vocabulary. The total entropy of the whole corpus is defined as .
@@ 4 Experiments
We evaluate the performance of TreeTok against the de-facto tokenizers such as BPE, WordPiece, and Unigram as primary baselines.
@@
For a fair comparison, we train all tokenizers from scratch on the lowercase version of the Wikitext-103 corpus McClosky et al. (2006) without any word boundary marker and set the same vocabulary size of 30,000. For BPE, WordPiece, and Unigram, we use the implementation and default training paradigm provided by the HuggingFace library333https://github.com/huggingface/tokenizers. Regarding the composition model, we train it with a context window of up to 512 characters. We use GPT2 implemented from HuggingFace444https://github.com/huggingface/transformers as our causal language model when computing the auto-regression loss. We present detailed configurations of our model and training setup in Appendix A.5.
@@
We compare our tokenizer with other tokenizers for morphological alignment using two datasets with gold-standard morphological segmentation. One is from the Morpho Challenge 2010 Workshop Kurimo et al. (2010) (Morpho), which contains 1,000 word forms with their segmentations corresponding to the surface forms of morpheme labels. The dataset contains instances of all kinds of morphological transformations, including inflection, derivation, and compounding. The other dataset is from Minixhofer et al. (2023) (Compound), which contains 759 compound words specifically designed to test the models’ capabilities in decompounding. We also use these morphological segmentation datasets to evaluate the induced morphological parse trees.
In addition, we evaluate the tokenizers using statistical metrics that have been shown to strongly correlate with the performance on downstream tasks. These metrics are calculated on the validation set of Wikitext-103.
@@ 4.1 Tokenization Quality
Baselines.
In addition to BPE, WordPiece, and Unigram, we also include two linguistically-motivated segmentation methods SECOS Riedl and Biemann (2016) and Morfessor Creutz and Lagus (2002) for reference.
@@
The superiority of Morfessor as shown in the table mainly comes from its much larger vocabulary. On the Compound task, we observe many cases where TreeTok segments a low-frequency constituent word into smaller pieces due to its absence in vocabulary, whereas Morfessor can find it in its extensive vocabulary. Because of the huge difference in vocabulary sizes, Morfessor is not directly comparable to the other tokenizers.
@@
Rényi | PPL | avg. #tokens | |
BPE | 44.66 | 107.76 | 26.58 |
WordPiece | 44.54 | 110.97 | 26.60 |
Unigram | 45.07 | 106.91 | 31.68 |
TreeTok | 44.82 | 107.26 | 25.99 |
@@ 4.2 Tree Structure Quality
For baseline composition models, we include Fast-R2D2 Hu et al. (2022), which is a variant of DIORA Drozdov et al. (2019), and an efficient variant of neural PCFG Yang et al. (2022).
@@
Morpho | Compound | |
EN. | EN. | |
Fast R2D2 | 67.69 | 48.96 |
Neural PCFG | 39.87 | 58.33 |
TreeTok | 90.10 | 86.20 |
w/o context | 70.00 | 63.02 |
w/o MorphOverriding | 75.99 | 46.35 |
w/o span loss | 86.79 | 73.70 |
@@
Figure 3 shows that the optimal size of an external vocabulary should be neither too large nor too small. According to our hypothesis that the compositional representation of subcomponents of a morpheme should be overridden by a high-level representation, ideally, the external vocabulary should contain all morphemes and only morphemes, because our model will trigger the soft morpheme overriding mechanism for every span that hits the external vocabulary. If we use BPE to construct this external vocabulary and the vocabulary size is too small, many morphemes, particularly those in the form of long standalone words, will not have the chance to be included in the vocabulary. Conversely, if the vocabulary size is too large, BPE will continue to merge across morphemes, creating longer spans that are not the smallest meaning-bearing units.
@@ 4.3 Case Studies
To complement our quantitative evaluations of different tokenizers and further examine their difference, we list some of their tokenizations over the same words in Table 4. Since our tokenizer’s performance is related to the tree structures learned by our composition model, we also include some of these tree structures in Figure 4.
Tokens produced by Unigram often include many characters. We observe that in addition to words that are out-of-domain (e.g., Japanese Romaji), single-character tokens mostly happen on short words with lengths of two or three (e.g., as, it, to, oil, etc.). This observation is consistent with notably longer tokenized sequences of Unigram observed in Table 2 compared to other tokenization methods. Note the overall likelihood of segmentation would be in a unigram language model. Therefore, we speculate that Unigram tends to split short words into characters since many characters have a dominantly high probability, while the length penalty resulting from the product of probabilities in short words is not significant.
BPE and WordPiece often violate morpheme boundaries and tokenize words into some intermediate “junk” tokens stored in their vocabulary during the bottom-up vocabulary construction. In Table 4, we can see many such examples such as fing, cles, and eness.
TreeTok aligns significantly better with morphology. Our vocabulary construction algorithm that brings the best from both BPE construction and Unigram pruning, helps eliminate “junk” tokens. In addition, with top-down matching under linguistic structure constraints, we alleviate the issue of breaking morpheme boundaries and excessively fragmenting words.
From Figure 4, we find the high-level structures of the trees learned by our model are generally accurate, whereas some of the low-level structures still do not make sense. This is because MorphOverriding allows the model to learn the most reasonable segmentation among different high-level splits based on context, while making the low-level structures less important, resulting in a certain degree of randomness within them.

### v2

# copyrightbox Unsupervised Morphological Tree Tokenizer

@@ Abstract
As a cornerstone in language modeling, tokenization involves segmenting text inputs into pre-defined atomic units. Conventional statistical tokenizers often disrupt constituent boundaries within words, thereby corrupting semantic information. To address this drawback, we introduce morphological structure guidance to tokenization and propose a deep model to induce character-level structures of words. Specifically, the deep model jointly encodes internal structures and representations of words with a mechanism named MorphOverriding to ensure the indecomposability of morphemes. By training the model with self-supervised objectives, our method is capable of inducing character-level structures that align with morphological rules without annotated training data. Based on the induced structures, our algorithm tokenizes words through vocabulary matching in a top-down manner. Empirical results indicate that the proposed method effectively retains complete morphemes and outperforms widely adopted methods such as BPE and WordPiece on both morphological segmentation tasks and language modeling tasks. Code is available at https://github.com/martianmartina/TreeTokenizer111Part of the work was done during Qingyang’s internship at Ant Group..
@@ 1 Introduction
According to linguistic theories, both words and sentences are believed to have internal structures Selkirk (1982); Marvin (2002); Cotterell and Schütze (2015). While sentence-level grammar induction methods based on deep neural networks are highly effective, whether these methods can be applied equally well to words remains under-explored. In this work, we systematically evaluate neural grammar induction methods at the word level, propose a hypothesis explaining their suboptimal performance, and empirically validate this hypothesis. Building on these insights, we introduce the first effective unsupervised neural grammar induction model at the word level and present a more morphologically aligned tokenizer that leverages our model as shown in Figure 1(bottom).
Our approach draws inspiration from syntactic composition models Maillard et al. (2017), where a sentence is encoded as a weighted sum over all composed root representations of its underlying binary parse trees via dynamic programming. Instead of composing a sentence from words, we apply composition models on characters in a word to induce its morphological parse tree. To train the composition model, we propose two self-supervised objectives akin to next token prediction and span prediction that effectively leverage both contextual information at the sentence level and semantic information at the subword level. Thus the model can learn to assign higher probabilities to morphological constituents of a word and induce the underlying morphological parse tree.
@@
In our experiments, which mainly focus on English, we train TreeTok and baselines on the Wikitext-103 corpus McClosky et al. (2006) and assess their performance on morphological segmentation tasks and language modeling tasks. Evaluation results indicate that TreeTok consistently outperforms BPE and WordPiece across all the tasks.
@@
We conduct empirical study on character-level neural parsing, identifying its limitations and proposing a novel explanation—lack of MorphOverriding—to account for its suboptimal performance.
Building on the MorphOverriding hypothesis, we introduce the first effective unsupervised neural model for character-level structure induction, addressing a critical gap in the field.
We show that our character-level structure induction method can be integrated into mainstream tokenizers to significantly enhance their performance on morphological tasks.
@@ 2 Related Work
Subword tokenization, with typical methods such as BPE Sennrich et al. (2016) and WordPiece Schuster and Nakajima (2012), has become customary in most NLP fields. BPE builds its vocabulary by repeatedly merging the most frequent subword unit pairs, whereas WordPiece selects pairs using the highest mutual information. During tokenization, BPE applies learned merge operations in the same order to new text initialized with characters while WordPiece iteratively finds the longest match in the vocabulary. Unigram Kudo (2018), another popular tokenizer, builds its vocabulary in the opposite direction: it starts with a large set of potential subwords and prunes them based on delta entropy in a unigram language model. Our tokenizer aims to build upon the advantages of these effective statistical tokenizers and augment them with unsupervised induced tree structures.
@@
In the line of work on unsupervised morphological segmentation, the most well-known model is Morfessor Creutz and Lagus (2002), along with its multiple variants Creutz and Lagus (2005); Grönroos et al. (2014, 2020). In Morfessor, an online search algorithm is utilized to apply a hierarchical word splitting strategy with a Minimum Description Length (MDL) Rissanen (1989) cost function. However, its lack of explicit control over vocabulary size makes it unsuitable for use as a tokenizer. In addition, although some studies Ataman and Federico (2018); Hou et al. (2023) find morphologically motivated segmentation can improve data-driven tokenizers, most other studies Machácek et al. (2018); Domingo et al. (2019); Sälevä and Lignos (2021) find no reliable improvement of such methods over BPE. According to Gallé (2019), the effectiveness of BPE lies in its superior compression capability. A more detailed discussion can be found in Mielke et al. (2021). Some other studies try to model morphological structures using Bayesian PCFGs Johnson et al. (2007) or a non-parametric Bayesian generalization of PCFGs Johnson et al. (2006). However, they are pure statistical models and do not utilize modern neural methodologies. Our method differs from previous unsupervised morphological methods in our character-based structures, thereby possessing the superior compression capability of BPE. Meanwhile, our method leverages modern neural methodologies to better utilize contextual and intra-word semantic information.
@@
In this work, we utilize a composition model to induce morphological structures. Composition models jointly learn representations and structures of a symbol sequence by transforming text encoding into a combinatorial optimization problem. Maillard et al. (2017) proposes a CKY-like Cocke (1969); Kasami (1966); Younger (1967) encoder, in which each constituent is represented as a weighted average of the set of composed representations computed from different splits of the constituent. Drozdov et al. (2019) proposes a deep inside-outside encoder Baker (1979); Lari and Young (1990), enabling the encoder to learn underlying structures via an auto-encoding objective. Recently, a series of studies Hu et al. (2024a, b) have been conducted to reduce the deep inside-outside encoder complexity from cubic to linear, on which our work is based.
@@ 3 Methodology
To tokenize a word where is the -th character, we aim to parse it into a binary tree and then tokenize it via top-down vocabulary matching. The parser is a deep composition model capable of jointly modeling the internal structures and representations of words. It is trained on sequences of words sampled from the corpus and optimized with self-supervised objectives that capture both intra-word compositionality and inter-word contextual dependencies. In the following sections, we sequentially introduce the composition model, training objectives, and the tree-based tokenization algorithm.
@@ 3.2 Training Objectives
The composition model is trained on contiguous sequences of words sampled from the corpus rather than isolated words. The overall loss for training the composition model is the summation of an intra-word auto-encoding loss and an inter-word auto-regression loss . The auto-encoding loss is based on predicting each character or morpheme from the rest of a word, leveraging intra-word structure. The auto-regression loss is based on predicting next word in the sequence that leverages contextual cues to disambiguate different underlying structures of a word. Under these objectives, the composition model learns to assign proper scores to each split point of a subword, benefiting from both intra-word compositionality and inter-word context.
Auto-encoding Loss (intra-word).
Auto-encoding is a common practice of training a composition model. For our character-level composition model, we try to predict each character based on its neighboring context representations and Hu et al. (2021). However, the auto-encoding objective turns out to be empirically ineffective when training our model probably because unlike word-level auto-encoding that requires selecting from tens of thousands of words in a vocabulary, here we only need to select from tens of characters, which is much less challenging.
To enhance learning efficacy, we propose predicting both individual characters and morphemes in the vocabulary . For instance, given the word windsurf, we mask out wind and let the model uncover the masked morpheme based on the visible part _surf. Analogous to the inside pass, the outside pass computes each outside representation in a top-down manner based on context information outside span , whose details are described in Appendix A.3. we use to predict each subword that belongs to :
where is the total number of subwords belonging to the vocabulary. 222Subword spans in the vocabulary may overlap (e.g., “asking” contains both “ask” and “king”), resulting in multiple competing candidates for prediction. Among the overlapping candidates, intuitively, it should be more reasonable to predict a constituent than other subword candidates from the context. Therefore, we assign a constituency weight to each subword in the objective, as detailed in Appendix A.4.
Auto-regression Loss (inter-word).
@@ 3.3 Tokenization
Given a constructed vocabulary, whose details are described later, we parse each word into a morphological tree and segment it via a top-down matching approach, as illustrated in Figure 1(bottom). Specifically, during the top-down traversal of a parse tree, we retain a subword and backtrack if the subword matches an entry in the vocabulary. Note that unsupervised structural learning is often imperfect, causing erroneous tokenization. For instance, an incorrect parse tree [[[book]e]d] may yield tokens book e d where e d should be merged. To address this issue, we propose a post-processing step to deal with mergeable pairs of segmented tokens. Specifically, we define the empirical probability of token as , where is the frequency of in the entire corpus and . Therefore, the probability of a certain merge is the production of the probabilities of all tokens. We find the optimal merge by searching for the one with maximum probability among all potential merges via dynamic programming. Detailed pseudo-code can be found in Appendix A.1.
@@
Tokenize the corpus via the top-down matching according to the current vocabulary. The total entropy of the whole corpus is defined as where .
@@ 4 Experiments
We focus on English in most of the experiments, but we also evaluated the composition model on Chinese in 4.2 and evaluated the tokenizer on German, which is a morphologically richer language than English, in a machine translation experiment in 4.1. We evaluate the performance of TreeTok against the de-facto tokenizers such as BPE, WordPiece, and Unigram as primary baselines.
@@
For a fair comparison, we train all tokenizers from scratch on the lowercase version of the WikiText-103 corpus Merity et al. (2017) without any word boundary marker and set the same vocabulary size of 30,000. For BPE, WordPiece, and Unigram, we use the implementation and default training paradigm provided by the HuggingFace library333https://github.com/huggingface/tokenizers. Regarding the composition model, we train it with a context window of up to 512 characters. We use GPT2 implemented from HuggingFace444https://github.com/huggingface/transformers as our causal language model when computing the auto-regression loss. We present detailed configurations of our model and training setup in Appendix A.5.
@@
We compare our tokenizer with other tokenizers for morphological alignment (detailed in 4.1) using two datasets with gold-standard morphological segmentation. One is from the Morpho Challenge 2010 Workshop Kurimo et al. (2010) (Morpho), which contains 1,000 word forms with their segmentations corresponding to the surface forms of morpheme labels. The dataset contains instances of all kinds of morphological transformations, including inflection, derivation, and compounding. The other dataset is from Minixhofer et al. (2023) (Compound), which contains 759 compound words specifically designed to test the models’ capabilities in decompounding. We also use these morphological segmentation datasets to evaluate the induced morphological parse trees (detailed in 4.2).
In addition, we evaluate the tokenizers using statistical metrics that have been shown to strongly correlate with the performance on downstream tasks (see Table 2). These metrics are calculated on the validation set of WikiText-103.
@@ 4.1 Tokenization Quality
| Rényi | PPL | BLEU | avg. #tokens | |
| BPE | 44.66 | 107.76 | 26.55 | 26.58 |
| WordPiece | 44.54 | 110.97 | - | 26.60 |
| Unigram | 45.07 | 106.91 | - | 31.68 |
| TreeTok | 44.82 | 107.26 | 26.68 | 25.99 |
@@
Machine Translation.
We conduct experiments on machine translation as a complementary. We use the fairseq framework 555https://github.com/facebookresearch/fairseq/blob/main/examples/translation/README.md##wmt14-english-to-german-convolutional to train a Transformer on WMT14 English to German from scratch and measure the performance by calculating the BLEU score on the official test split.
We compare the model’s performance when the tokenizer is BPE and TreeTok, respectively. We use the same model training setups. For the two tokenizers, the vocabulary size and basic characters are exactly the same. The results from Table 2 show that TreeTok is slightly better than BPE. Based on the results, TreeTok can improve alignment with morphology on top of BPE, while the new segmentation does not compromise downstream task performance.
@@ 4.2 Tree Structure Quality
For baseline composition models, we include Fast-R2D2 Hu et al. (2022), which is a variant of DIORA Drozdov et al. (2019), and an efficient variant of neural PCFG Yang et al. (2022). We choose PCFG and R2D2 as our baselines because they represent two different classic approaches to modeling composition: PCFGs are based on underlying grammar structures, while R2D2 (a neural inside algorithm) searches for the optimal information compression structure through binary composition of tokens.
@@
| Morpho | Compound | Word Seg. | |
| EN. | EN. | ZH. | |
| Fast R2D2 | 67.69 | 48.96 | — |
| Neural PCFG | 39.87 | 58.33 | 74.26 |
| TreeTok | 90.10 | 86.20 | — |
| w/o context | 70.00 | 63.02 | — |
| w/o MorphOverriding | 75.99 | 46.35 | 99.24 |
| w/o span weights | 89.42 | 78.39 | — |
| w/o span loss | 86.79 | 73.70 | — |
@@
In addition, we train both Neural-PCFG and our composition model on Chinese wiki Xu and Lapata (2019) and evaluate the recall of word boundary against Penn Chinese Treebank XUE et al. (2005). Our model can achieve a word boundary recall of 99.24% without MorphOverriding. Chinese is a language system that can be considered simplistic in terms of its internal structure of words where most of the time, each Chinese character (referred to as a hàn zì) represents one morpheme, and there are always explicit boundaries between morphemes without orthographic changes during word formation from characters. These features make compositionality applicable in most cases, thus alleviating the difficulty of modelling intra-word structures. Hence, comparing with the poor performance of w/o MorphOverriding on the English dataset (Compound), we can conclude that the difficulties of modelling the internal word structure vary greatly across languages, and MorphOverriding is effective and necessary for languages with more challenging morphology structures.
@@
Figure 3 shows that the optimal size of an external vocabulary should be neither too large nor too small. According to our hypothesis that the compositional representation of subcomponents of a morpheme should be overridden by a high-level representation, ideally, the external vocabulary should contain all morphemes and only morphemes, because our model will trigger the soft morpheme overriding mechanism for every span that hits the external vocabulary. If BPE is used and the vocabulary is too small, many morphemes (especially longer standalone words) are excluded. Conversely, if it is too large, BPE merges across morphemes, creating spans larger than the smallest meaning-bearing units.
@@ 4.3 Case Studies
To further examine the difference between tokenizers, we list their tokenizations in Table 4 and tree structures induced by our composition model in Figure 4.
Tokens produced by Unigram often include many characters. BPE and WordPiece often violate morpheme boundaries and tokenize words into some intermediate “junk” tokens introduced during the bottom-up vocabulary construction, such as fing, cles, and eness in Table 4.
TreeTok aligns significantly better with morphology. By merging the best of BPE and Unigram pruning, our vocabulary construction algorithm eliminates “junk” tokens. Meanwhile, top-down matching under linguistic constraints prevents excessive word fragmentation and morpheme boundary breaks.
In Figure 4, our model’s high-level tree structures are generally accurate, although some low-level structures appear random, since MorphOverriding prioritizes the most reasonable high-level segmentations based on context, making lower-level details less important.